OpenStudy (anonymous):

Solve: −4(x + 1) − 3 = −3(x − 4)

OpenStudy (anonymous):

-4x-4-3 = -3x+12 Rearrange; -3x+12=-4x-4-3 like terms together; -3x+4x = -4-3-12 x = ????

OpenStudy (anonymous):

-4x-4-3=-3x+12 -4x-7=-3x+12

OpenStudy (neo92):

didnt u get it yet?

OpenStudy (anonymous):

1) distribute everywhere possible 2) after distributing you get -4x-4-3=-3x+12 3) now combine like terms to get -4x-7=-3x+12 4) add 3x on both sides to get -1x-7=12 5) add 7 to both sides to get -1x=19 6) divide both sides by -1 to get x=-19 Hope that helps :D